Plan for PURE Math 2012 Seminar

Week 3

Monday: Distance coordinates and configuration spaces Tuesday: Central configurations in the 3- and n-body problems Wednesday: Thursday:

PURE Math 2012 - Seminar Week 3 Discussions - Background on Central Configurations

Day 1: Distance coordinates and configuration spaces

Background

We have seen in lecture that one way to describe configurations of n points p_{1}, …, p_{n} in a Euclidean space up to Euclidean motions (translations and rotations) is to introduce the mutual distances r_{ij} = ∥p_{i} − p_{j}∥ for 1 ≤ i < j ≤ n as coordinates.